OpenStudy (yash2651995):

i'm not able to understand this: (this is WRT ODE course.) Linear - When do we say that F(x; y; y1; : : : ; y(n)) = 0 linear? Write F(x; y; y1; : : : ; y(n)) = 0 as L(y)(x) = L(y; y1; : : : ; y(n)) = f(x); and now L is a linear transformation from Cn = the space of functions which are at least n-times di erentiable to F = the space of functions. If these functions are de ned on an open set  R, we say that the ODE is de ned on . (Why do we take an open set?)

OpenStudy (anonymous):

"dened" ?

OpenStudy (yash2651995):

defined* both places sorry about typo

OpenStudy (anonymous):

I believe the "open set" requirement is there for the same reason a function \(f:\mathbb{R}\to\mathbb{R}\) can only be differentiable over an open interval \((a,b)\subset\mathbb{R}\).